OpenStudy (anonymous):

Integrate x(x^2+1)^-1/2

OpenStudy (anonymous):

\[\frac{ x }{ \sqrt{x^2+1} }\]

OpenStudy (anonymous):

Maybe if you post this in the Math section, you'll find some answers. Have a nice day! ^-^

OpenStudy (anonymous):

u would get better help if u posted this on the math section to get to the math section go to the top left where it says Find more subjects then go to math and post it well have a good day ^-^

OpenStudy (anonymous):

I was in the wrong subject... I got my answer when I posted in the math section. Thank you guys!

OpenStudy (anonymous):

np

OpenStudy (anonymous):

what was the answer i am curious

OpenStudy (anonymous):

\[\sqrt{x^2+1}+C\] would be the anti-derivative.

OpenStudy (anonymous):

The process was a lot more exciting lol

OpenStudy (anonymous):

lol